find and fill the missing items.


2020, first half 2020, first half


Gross purchases P350,000


Purchase returns and discounts 8% of the overall purchases value


Beginning inventory P82,000


Ending inventory 30% higher than the beginning inventory


Net sales P910,000


Distribution expenses 25% of cost of goods sold


Administrative expenses 15% of gross profit


Income tax expense 20% of net income before taxes


Net income ? (1)


Cost of goods sold ? (2)


Gross profit ? (3)


Distribution expenses ? (4)


Income tax expenses ? (5)



1
Expert's answer
2021-09-20T16:39:51-0400

Solution:

Cost of goods sold = Beginning inventory + net purchases – ending inventory

Beginning inventory = 82,000

Gross purchases = 350,000

Purchases returns & discounts = 350,000"\times 8\%" = 28,000

Net purchases = Purchases – purchases returns & discounts = 350,000 – 28,000 = 322,000

Ending inventory = 82,000"\times 130\%" = 106,600

Cost of goods sold = 82,000 + 322,000 – 106,600 = 297,400

Gross profit = Net sales – Cost of goods sold

Net sales = 910,000

Gross profit = 910,000 – 297,400 = 612,600

Distribution expenses = 25"\%" of COGS = 297,400 "\times 25\%" = 74,350

Administrative expenses = 15"\%" of Gross profit = 612,600 "\times 15\%" = 91,890

 

Net income before taxes = Gross profit – distribution expenses – administrative expenses

= 612,600 – 74,350 – 91,890 = 446,360

Net income before taxes = 446,360

Income tax expenses = 20"\%" of net income before taxes = 446,360 "\times 20\%" = 89,272

Net income = Net income before taxes – Income tax expenses = 446,360 – 89,272 = 357,088


Net income (1) = P357,088

 

Cost of goods sold (2) = P297,400

 

Gross profit (3) = P612,600

 

Distribution expenses (4) = P74,350

 

Income tax expenses (5) = P89,272